# All-in-One: Heterogeneous Interaction Modeling for Cold-Start Rating Prediction

### v2

@@ 17       Update model parameters by gradient descent;
TABLE V: Overall Performance in Three Cold-Start Scenarios on $\mathsf{Douban}$
@@
UC NeuMF $0.4443_{\small(.0160)}$ $0.3334_{\small(.0199)}$ $0.4056_{\small(.0164)}$ $0.5872_{\small(.0124)}$ $0.3471_{\small(.0191)}$ $0.5623_{\small(.0130)}$ $0.7261_{\small(.0071)}$ $0.3765_{\small(.0168)}$ $0.7113_{\small(.0080)}$
Wide&Deep $0.5442_{\small(.0105)}$ $0.7725_{\small(.0083)}$ $0.4443_{\small(.0175)}$ $0.5566_{\small(.0031)}$ $0.7595_{\small(.0081)}$ $0.4501_{\small(.0120)}$ $0.5826_{\small(.0032)}$ $0.7376_{\small(.0070)}$ $0.4783_{\small(.0068)}$
DeepFM $0.5133_{\small(.0178)}$ $0.7261_{\small(.0242)}$ $0.4141_{\small(.0250)}$ $0.5133_{\small(.0239)}$ $0.7016_{\small(.0293)}$ $0.4115_{\small(.0280)}$ $0.5281_{\small(.0147)}$ $0.6707_{\small(.0286)}$ $0.4312_{\small(.0193)}$
AFN $0.5918_{\small(.0119)}$ $0.8041_{\small(.0171)}$ $0.4919_{\small(.0198)}$ $0.6088_{\small(.0083)}$ $0.7981_{\small(.0182)}$ $0.4992_{\small(.0157)}$ $0.6442_{\small(.0108)}$ $0.7875_{\small(.0198)}$ $0.5407_{\small(.0038)}$
GraphRec $0.6065_{\small(.0279)}$ $0.5073_{\small(.0366)}$ $0.5477_{\small(.0326)}$ $\mathbf{0.7549_{\small(.0238)}}$ $0.5605_{\small(.0373)}$ $\mathbf{0.7220_{\small(.0267)}}$ $\mathbf{0.8771_{\small(.0172)}}$ $0.6206_{\small(.0356)}$ $\mathbf{0.8631_{\small(.0189)}}$
MAMO $0.6098_{\small(.0007)}$ $0.7356_{\small(.0005)}$ $0.5101_{\small(.0005)}$ $0.6796_{\small(.0005)}$ $0.7594_{\small(.0002)}$ $0.5741_{\small(.0004)}$ $0.8006_{\small(.0003)}$ $0.8060_{\small(.0002)}$ $0.7209_{\small(.0004)}$
TaNP $0.6408_{\small(.0340)}$ $0.9020_{\small(.0104)}$ $0.5465_{\small(.0447)}$ $0.6906_{\small(.0308)}$ $0.9100_{\small(.0094)}$ $0.5919_{\small(.0422)}$ $0.7824_{\small(.0386)}$ $0.9259_{\small(.0104)}$ $0.7044_{\small(.0530)}$
MeLU $0.4542_{\small(.0480)}$ $0.6452_{\small(.0181)}$ $0.3463_{\small(.0457)}$ $0.5054_{\small(.0706)}$ $0.6623_{\small(.0269)}$ $0.3815_{\small(.0693)}$ $0.5953_{\small(.0798)}$ $0.6930_{\small(.0352)}$ $0.4723_{\small(.0885)}$
HIRE $\mathbf{0.7152_{\small(.0391)}}$ $\mathbf{0.9269_{\small(.0131)}}$ $\mathbf{0.6595_{\small(.0574)}}$ $0.7459_{\small(.0248)}$ $\mathbf{0.9328_{\small(.0122)}}$ $0.6837_{\small(.0420)}$ $0.8046_{\small(.0130)}$ $\mathbf{0.9479_{\small(.0111)}}$ $0.7521_{\small(.0108)}$
IC NeuMF $0.3919_{\small(.0442)}$ $0.4305_{\small(.0428)}$ $0.3050_{\small(.0183)}$ $0.4688_{\small(.0167)}$ $0.4423_{\small(.0413)}$ $0.3693_{\small(.0212)}$ $0.5604_{\small(.0088)}$ $0.4644_{\small(.0395)}$ $0.4536_{\small(.0309)}$
Wide&Deep $0.2285_{\small(.0080)}$ $0.4496_{\small(.0173)}$ $0.1787_{\small(.0059)}$ $0.2500_{\small(.0153)}$ $0.4453_{\small(.0170)}$ $0.1878_{\small(.0127)}$ $0.2937_{\small(.0130)}$ $0.4533_{\small(.0172)}$ $0.2168_{\small(.0105)}$
DeepFM $0.2390_{\small(.0023)}$ $0.4723_{\small(.0155)}$ $0.1856_{\small(.0029)}$ $0.2574_{\small(.0019)}$ $0.4665_{\small(.0151)}$ $0.1881_{\small(.0053)}$ $0.2970_{\small(.0068)}$ $0.4731_{\small(.0148)}$ $0.2130_{\small(.0074)}$
AFN $0.2600_{\small(.0271)}$ $0.5014_{\small(.0255)}$ $0.2044_{\small(.0217)}$ $0.2875_{\small(.0210)}$ $0.4993_{\small(.0238)}$ $0.2160_{\small(.0175)}$ $0.3298_{\small(.0259)}$ $0.5056_{\small(.0265)}$ $0.2405_{\small(.0183)}$
GraphRec $0.3460_{\small(.0298)}$ $0.3973_{\small(.0360)}$ $0.2847_{\small(.0393)}$ $0.4328_{\small(.0328)}$ $0.4273_{\small(.0409)}$ $0.3597_{\small(.0436)}$ $0.5388_{\small(.0270)}$ $0.4639_{\small(.0424)}$ $0.4508_{\small(.0469)}$
MAMO $0.5980_{\small(.0011)}$ $0.7250_{\small(.0005)}$ $0.4986_{\small(.0010)}$ $\mathbf{0.6696_{\small(.0011)}}$ $0.7495_{\small(.0004)}$ $0.5639_{\small(.0012)}$ $\mathbf{0.7943_{\small(.0006)}}$ $0.7978_{\small(.0003)}$ $\mathbf{0.7120_{\small(.0006)}}$
TaNP $0.4945_{\small(.0303)}$ $0.8502_{\small(.0090)}$ $0.3808_{\small(.0326)}$ $0.5479_{\small(.0396)}$ $0.8590_{\small(.0100)}$ $0.4230_{\small(.0462)}$ $0.6328_{\small(.0595)}$ $0.8740_{\small(.0135)}$ $0.5107_{\small(.0750)}$
MeLU $0.5087_{\small(.0363)}$ $0.6650_{\small(.0221)}$ $0.3876_{\small(.0335)}$ $0.5619_{\small(.0685)}$ $0.6859_{\small(.0371)}$ $0.4354_{\small(.0822)}$ $0.6512_{\small(.0902)}$ $0.7179_{\small(.0493)}$ $0.5770_{\small(.0576)}$
HIRE $\mathbf{0.6128_{\small(.0467)}}$ $\mathbf{0.8926_{\small(.0159)}}$ $\mathbf{0.5496_{\small(.0679)}}$ $0.6463_{\small(.0307)}$ $\mathbf{0.8983_{\small(.0147)}}$ $\mathbf{0.5708_{\small(.0513)}}$ $0.7066_{\small(.0067)}$ $\mathbf{0.9122_{\small(.0136)}}$ $0.6305_{\small(.0196)}$
U&I C NeuMF $0.2763_{\small(.0055)}$ $0.3898_{\small(.0098)}$ $0.2266_{\small(.0119)}$ $0.2857_{\small(.0130)}$ $0.3616_{\small(.0118)}$ $0.2237_{\small(.0099)}$ $0.3340_{\small(.0235)}$ $0.3597_{\small(.0191)}$ $0.2478_{\small(.0079)}$
Wide&Deep $0.0910_{\small(.0210)}$ $0.1615_{\small(.0143)}$ $0.0819_{\small(.0193)}$ $0.0820_{\small(.0196)}$ $0.1435_{\small(.0118)}$ $0.0731_{\small(.0144)}$ $0.0753_{\small(.0230)}$ $0.1303_{\small(.0153)}$ $0.0637_{\small(.0178)}$
DeepFM $0.0682_{\small(.0324)}$ $0.1433_{\small(.0742)}$ $0.0596_{\small(.0281)}$ $0.0596_{\small(.0330)}$ $0.1248_{\small(.0663)}$ $0.0521_{\small(.0292)}$ $0.0561_{\small(.0332)}$ $0.1106_{\small(.0610)}$ $0.0463_{\small(.0263)}$
AFN $0.0609_{\small(.0159)}$ $0.1484_{\small(.0195)}$ $0.0552_{\small(.0133)}$ $0.0588_{\small(.0167)}$ $0.1286_{\small(.0192)}$ $0.0527_{\small(.0126)}$ $0.0559_{\small(.0155)}$ $0.1123_{\small(.0182)}$ $0.0485_{\small(.0142)}$
GraphRec $0.3568_{\small(.0283)}$ $0.3900_{\small(.0168)}$ $0.2624_{\small(.0178)}$ $0.4066_{\small(.0271)}$ $0.4043_{\small(.0165)}$ $0.2885_{\small(.0234)}$ $0.4700_{\small(.0181)}$ $0.4276_{\small(.0128)}$ $0.3326_{\small(.0149)}$
MAMO $0.6009_{\small(.0012)}$ $0.7278_{\small(.0004)}$ $0.5037_{\small(.0011)}$ $0.6750_{\small(.0016)}$ $0.7531_{\small(.0005)}$ $0.5721_{\small(.0016)}$ $\mathbf{0.7976_{\small(.0006)}}$ $0.8007_{\small(.0003)}$ $\mathbf{0.7173_{\small(.0008)}}$
TaNP $0.6067_{\small(.0244)}$ $0.8734_{\small(.0040)}$ $0.4982_{\small(.0217)}$ $\mathbf{0.6962_{\small(.0089)}}$ $0.8903_{\small(.0030)}$ $0.5805_{\small(.0166)}$ $0.7755_{\small(.0174)}$ $0.9100_{\small(.0039)}$ $0.6690_{\small(.0171)}$
MeLU $0.5032_{\small(.0173)}$ $0.6737_{\small(.0070)}$ $0.3934_{\small(.0111)}$ $0.5544_{\small(.0273)}$ $0.6875_{\small(.0116)}$ $0.4355_{\small(.0228)}$ $0.6717_{\small(.0125)}$ $0.7294_{\small(.0078)}$ $0.5626_{\small(.0136)}$
HIRE $\mathbf{0.6266_{\small(.0195)}}$ $\mathbf{0.8902_{\small(.0214)}}$ $\mathbf{0.5416_{\small(.0586)}}$ $0.6953_{\small(.0169)}$ $\mathbf{0.9017_{\small(.0194)}}$ $\mathbf{0.6148_{\small(.0324)}}$ $0.7833_{\small(.0251)}$ $\mathbf{0.9247_{\small(.0166)}}$ $0.7106_{\small(.0200)}$
VI Experimental Studies
In this section, we give the test setting (§ VI-A) and report our substantial experimental results in the following facets: (1) Compare the effectiveness of HIRE with the state-of-the-art approaches under three cold-start scenarios (§ VI-B). (2) Compare the test efficiency of HIRE with the baseline approaches (§ VI-C). (3) Investigate the sensitivity of HIRE regarding key hyper-parameter configurations (§ VI-D). (4) Study the influence of different attention layers via an ablation study (§ VI-E). (5) Conduct a case study for HIRE on a movie rating task. (§ VI-F).
VI-A Experimental Setup
Datasets: We use 3 widely used datasets to evaluate HIRE and the baseline approaches, whose profiles are summarized in Table II. $\mathsf{MovieLens}$-$\mathsf{1M}$ [46] contains the ratings of users for movies. We randomly split 80% of users for training and 20% of users as the cold-start test users. Movies whose release years are before 1997 are used for training and the remaining are used for cold-start prediction. Both the users and movies are associated with rich attributes. $\mathsf{Douban}$ [47, 48] contains the ratings of users on music, which are extracted from a rating website Douban. Users and music are randomly split by 70% and 30% for training and cold-start testing, respectively. The dataset also provides friendship relations between users. Since users/items do not have attributes, we use the embeddings of the user/item ID as the attributes, which are generated by learnable linear weights. $\mathsf{Bookcrossing}$ [49] is collected from Book-Crossing, containing ratings of users on books. Similar to $\mathsf{Douban}$, both users and books are randomly split by 70% and 30% for training and cold-start testing, respectively.
Baselines: To comprehensively evaluate HIRE, we compare with the following $10$ baselines which fall into 4 categories, i.e., (1) $4$ neural CF-based approaches, NeuMF  [^1] https://github.com/hexiangnan/neural_collaborative_filtering, DeepFM  [^2] https://github.com/shenweichen/DeepCTR-Torch, Wide&Deep 22footnotemark: 2 and AFN 22footnotemark: 2, (2) a social recommendation approach, GraphRec  [^3] https://github.com/wenqifan03/GraphRec-WWW19, (3) $2$ HIN-based approaches, MetaHIN  [^4] https://github.com/rootlu/MetaHIN and GraphHINGE  [^5] https://github.com/Jinjiarui/GraphHINGE, (4) $3$ meta-learning approaches for cold-start recommendation, MAMO  [^6] https://github.com/dongmanqing/Code-for-MAMO, TaNP  [^7] https://github.com/IIEdm/TaNP and MeLU  [^8] https://github.com/hoyeoplee/MeLU.
  • CF-based baselines: NeuMF [8], Wide&Deep [25], DeepFM [26] and AFN [27] adopt different neural network layers to model the interactions from user-item pairs. The CF-based approaches are not specifically designed for cold-start.
  • Social recommendation: GraphRec [15] models the users and items respectively by aggregating their local interactions via Graph Neural Network. Aggregation of the social relation among users is used to enhance user embeddings. We only compare GraphRec on $\mathsf{Douban}$ where the social relationship between users is available.
  • HIN-based baselines: GraphHINGE [21] aggregates rich interactive patterns from an HIN. MetaHIN [33] incorporates multifaceted semantic contexts induced by meta-paths. We test MetaHIN and GraphHINGE only on $\mathsf{MovieLens}$-$\mathsf{1M}$ which provides sufficient attributes for constructing an HIN. Specifically, we define user, movie, genre, occupation, and age as different node types and create links between nodes, such as movie-genre interaction, user-age interaction, and user-movie interaction. For the other two datasets $\mathsf{Bookcrossing}$ and $\mathsf{Douban}$, due to lack of sufficient attributes of user/item, the two baselines cannot be applied to them.
  • Meta-learning baselines: MAMO [24], MeLU [23] and TaNP [22] adopt meta-learning for cold-start recommendation in a multi-task fashion. MAMO and MeLU support all the three cold-start scenarios while TaNP only supports user cold-start originally. Here, we extend the task settings of TaNP to item cold-start and user & item cold-start.
Implementation details: We introduce the model and training details of HIRE as well as the baselines. For HIRE, the model is equipped with $3$ HIM blocks in which each MHSA layer has $8$ heads and the hidden dimension of each head is set to $16$. The model is trained by a LAMB optimizer [50] with $\beta=(0.9,0.999)$ and $\epsilon=10^{-6}$, and a Lookahead [51] wrapper with slow update rate $\alpha=0.5$ and $k=6$ steps between updates. We use a flat-then-anneal learning rate scheduler which flats at the base learning rate for $70\%$ of steps, and then anneals following a cosine schedule to 00 by the end of training. We set the base learning rate to $10^{-3}$ and the gradient clip threshold to $1.0$. The optimizer and the scheduler are widely used for training language models composed of deep MHSA layers. Regarding the prediction contexts, we set the number of users and items in each context to 32 by default. In the training and test, $90\%$ of observed ratings are masked for prediction and the remaining $10\%$ are associated with the context input. The learning framework of HIRE is built on PyTorch [52].
For the baseline approaches, we use their released source code and keep the hyper-parameters as their default settings. All the models are trained until convergence. For the four CF-based approaches, GraphRec and GraphHINGE, to achieve a fair comparison, we use all the observed user-item ratings in the sampled training contexts, together with the $10\%$ unmasked user-item ratings in the test context as the ground truth to train the models. For meta-learning approaches, in each task, $10\%$ of the observed ratings are used as the ground truth in the support set and the remaining $90\%$ are used in the query set. Training and testing are conducted on one Tesla V100 with 16GB memory.
Evaluation Metrics: We adopt three commonly used metrics to evaluate the recommendation performance in our testing, including: Precision, Mean Average Precision (MAP) and Normalized Discounted Cumulative Gain [53] (NDCG). Top $k$ actual rating values sorted by predicted rating values are used to calculate the above metrics, where $k\in\{5,7,10\}$.

### v3

@@ VI-A Experimental Setup
TABLE V: Overall Performance in 3 Cold-Start Scenarios on $\mathsf{Douban}$ (%)
@@
UC Popularity $51.43_{\small(0.25)}$ $86.58_{\small(0.06)}$ $40.01_{\small(0.29)}$ $56.03_{\small(0.27)}$ $87.24_{\small(0.06)}$ $43.12_{\small(0.34)}$ $64.46_{\small(0.23)}$ $88.52_{\small(0.06)}$ $51.20_{\small(0.37)}$
TransGNN $38.54_{\small(0.90)}$ $38.45_{\small(1.35)}$ $21.76_{\small(1.24)}$ $38.69_{\small(2.10)}$ $38.51_{\small(1.53)}$ $23.55_{\small(1.51)}$ $38.96_{\small(1.30)}$ $38.64_{\small(1.80)}$ $25.60_{\small(2.10)}$
NeuMF $44.43_{\small(1.60)}$ $33.34_{\small(1.99)}$ $40.56_{\small(1.64)}$ $58.72_{\small(1.24)}$ $34.71_{\small(1.91)}$ $56.23_{\small(1.30)}$ $72.61_{\small(0.71)}$ $37.65_{\small(1.68)}$ $71.13_{\small(0.80)}$
Wide&Deep $54.42_{\small(1.05)}$ $73.76_{\small(0.70)}$ $44.43_{\small(1.75)}$ $55.66_{\small(0.31)}$ $75.95_{\small(0.81)}$ $45.01_{\small(1.20)}$ $58.26_{\small(0.32)}$ $77.25_{\small(0.83)}$ $47.83_{\small(0.68)}$
DeepFM $51.32_{\small(1.78)}$ $67.07_{\small(2.86)}$ $40.15_{\small(1.69)}$ $51.33_{\small(2.39)}$ $70.16_{\small(2.93)}$ $41.15_{\small(2.80)}$ $52.81_{\small(1.47)}$ $72.61_{\small(2.42)}$ $43.12_{\small(1.93)}$
AFN $59.18_{\small(1.19)}$ $78.75_{\small(1.98)}$ $49.19_{\small(1.98)}$ $60.88_{\small(0.83)}$ $79.81_{\small(1.82)}$ $49.92_{\small(1.57)}$ $64.42_{\small(1.08)}$ $80.41_{\small(1.71)}$ $54.07_{\small(0.38)}$
GraphRec $60.65_{\small(2.79)}$ $50.73_{\small(3.66)}$ $54.77_{\small(3.26)}$ $\mathbf{75.49_{\small(2.38)}}$ $56.05_{\small(3.73)}$ $\mathbf{72.20_{\small(2.67)}}$ $\mathbf{87.71_{\small(1.72)}}$ $62.06_{\small(3.56)}$ $\mathbf{86.31_{\small(1.89)}}$
MAMO $60.98_{\small(0.07)}$ $73.56_{\small(0.05)}$ $51.01_{\small(0.05)}$ $67.96_{\small(0.05)}$ $75.94_{\small(0.02)}$ $57.41_{\small(0.04)}$ $80.06_{\small(0.03)}$ $80.60_{\small(0.02)}$ $72.09_{\small(0.04)}$
TaNP $64.08_{\small(3.40)}$ $90.20_{\small(1.04)}$ $54.65_{\small(4.47)}$ $69.06_{\small(3.08)}$ $91.00_{\small(0.94)}$ $59.19_{\small(4.22)}$ $78.24_{\small(3.86)}$ $92.59_{\small(1.04)}$ $70.44_{\small(5.30)}$
MeLU $45.42_{\small(4.80)}$ $64.52_{\small(1.81)}$ $34.63_{\small(4.57)}$ $50.54_{\small(7.06)}$ $66.23_{\small(2.69)}$ $38.15_{\small(6.93)}$ $59.53_{\small(7.98)}$ $69.30_{\small(3.52)}$ $47.23_{\small(8.85)}$
HIRE $\mathbf{71.52_{\small(3.91)}}$ $\mathbf{92.69_{\small(1.31)}}$ $\mathbf{65.95_{\small(5.74)}}$ $74.59_{\small(2.48)}$ $\mathbf{93.28_{\small(1.22)}}$ $68.37_{\small(4.20)}$ $80.46_{\small(1.30)}$ $\mathbf{94.79_{\small(1.11)}}$ $75.21_{\small(1.08)}$
IC Popularity $47.35_{\small(0.40)}$ $84.35_{\small(0.22)}$ $36.10_{\small(0.45)}$ $54.63_{\small(0.08)}$ $85.44_{\small(0.12)}$ $42.06_{\small(0.30)}$ $64.85_{\small(0.09)}$ $87.12_{\small(0.09)}$ $51.76_{\small(0.14)}$
TransGNN $54.67_{\small(2.30)}$ $50.73_{\small(2.01)}$ $44.49_{\small(0.84)}$ $59.42_{\small(0.99)}$ $55.10_{\small(0.59)}$ $47.38_{\small(0.37)}$ $61.33_{\small(1.15)}$ $57.51_{\small(0.47)}$ $49.39_{\small(0.25)}$
NeuMF $39.19_{\small(4.42)}$ $43.05_{\small(4.28)}$ $30.50_{\small(1.83)}$ $46.88_{\small(1.67)}$ $44.23_{\small(4.13)}$ $36.93_{\small(2.12)}$ $56.04_{\small(0.88)}$ $46.44_{\small(3.95)}$ $45.36_{\small(3.09)}$
Wide&Deep $22.85_{\small(0.80)}$ $44.27_{\small(1.78)}$ $17.87_{\small(0.59)}$ $25.00_{\small(1.53)}$ $44.53_{\small(1.70)}$ $18.78_{\small(1.27)}$ $29.37_{\small(1.30)}$ $45.33_{\small(1.72)}$ $21.68_{\small(1.05)}$
DeepFM $23.90_{\small(0.23)}$ $46.33_{\small(0.21)}$ $18.56_{\small(0.29)}$ $25.74_{\small(0.19)}$ $46.65_{\small(1.51)}$ $18.81_{\small(0.53)}$ $29.70_{\small(0.68)}$ $47.31_{\small(1.48)}$ $21.30_{\small(0.74)}$
AFN $26.00_{\small(2.71)}$ $49.15_{\small(2.67)}$ $20.44_{\small(2.17)}$ $28.75_{\small(2.10)}$ $49.93_{\small(2.38)}$ $21.60_{\small(1.75)}$ $32.98_{\small(2.59)}$ $50.56_{\small(2.65)}$ $24.05_{\small(1.83)}$
GraphRec $34.60_{\small(2.98)}$ $39.73_{\small(3.60)}$ $28.47_{\small(3.93)}$ $43.28_{\small(3.28)}$ $42.73_{\small(4.09)}$ $35.97_{\small(4.36)}$ $53.88_{\small(2.70)}$ $46.39_{\small(4.24)}$ $45.08_{\small(4.69)}$
MAMO $59.80_{\small(0.11)}$ $72.50_{\small(0.05)}$ $49.86_{\small(0.10)}$ $\mathbf{66.96_{\small(0.11)}}$ $74.95_{\small(0.04)}$ $56.39_{\small(0.12)}$ $\mathbf{79.43_{\small(0.06)}}$ $79.78_{\small(0.03)}$ $\mathbf{71.20_{\small(0.06)}}$
TaNP $49.45_{\small(3.03)}$ $85.02_{\small(0.90)}$ $38.08_{\small(3.26)}$ $54.79_{\small(3.96)}$ $85.90_{\small(1.00)}$ $42.30_{\small(4.62)}$ $63.28_{\small(5.95)}$ $87.40_{\small(1.35)}$ $51.07_{\small(7.50)}$
MeLU $50.87_{\small(3.63)}$ $66.50_{\small(2.21)}$ $38.76_{\small(3.35)}$ $56.19_{\small(6.85)}$ $68.59_{\small(3.71)}$ $43.54_{\small(8.22)}$ $65.12_{\small(9.02)}$ $71.79_{\small(4.93)}$ $57.70_{\small(5.76)}$
HIRE $\mathbf{61.28_{\small(4.67)}}$ $\mathbf{89.26_{\small(1.59)}}$ $\mathbf{54.96_{\small(6.79)}}$ $64.63_{\small(3.07)}$ $\mathbf{89.83_{\small(1.47)}}$ $\mathbf{57.08_{\small(5.13)}}$ $70.66_{\small(0.67)}$ $\mathbf{91.22_{\small(1.36)}}$ $63.05_{\small(1.96)}$
U&I C Popularity ${47.54_{\small(0.39)}}$ $87.82_{\small(0.19)}$ ${42.24_{\small(0.57)}}$ $54.10_{\small(0.26)}$ $87.95_{\small(0.05)}$ ${46.47_{\small(0.27)}}$ $70.82_{\small(0.05)}$ ${90.14_{\small(0.06)}}$ $62.91_{\small(1.12)}$
TransGNN $27.33_{\small(1.15)}$ $30.34_{\small(1.50)}$ $25.07_{\small(0.71)}$ $30.47_{\small(1.65)}$ $32.91_{\small(1.88)}$ $27.33_{\small(1.01)}$ $30.67_{\small(2.30)}$ $33.62_{\small(2.26)}$ $29.11_{\small(1.02)}$
NeuMF $27.63_{\small(0.55)}$ $35.97_{\small(1.91)}$ $22.02_{\small(0.63)}$ $28.57_{\small(1.30)}$ $36.16_{\small(1.18)}$ $22.37_{\small(0.99)}$ $33.40_{\small(2.35)}$ $38.98_{\small(0.98)}$ $24.78_{\small(0.79)}$
Wide&Deep $8.42_{\small(2.72)}$ $13.03_{\small(1.53)}$ $7.30_{\small(1.67)}$ $8.48_{\small(2.56)}$ $14.35_{\small(1.18)}$ $7.31_{\small(1.44)}$ $8.57_{\small(2.02)}$ $16.15_{\small(1.43)}$ $8.32_{\small(1.77)}$
DeepFM $5.78_{\small(3.82)}$ $11.03_{\small(6.67)}$ $4.82_{\small(2.83)}$ $5.96_{\small(3.30)}$ $12.48_{\small(6.63)}$ $5.21_{\small(2.92)}$ $7.52_{\small(0.36)}$ $14.50_{\small(1.82)}$ $6.12_{\small(0.84)}$
AFN $5.26_{\small(0.98)}$ $11.23_{\small(1.82)}$ $4.76_{\small(0.26)}$ $5.88_{\small(1.67)}$ $12.86_{\small(1.92)}$ $5.27_{\small(1.26)}$ $6.39_{\small(1.00)}$ $14.84_{\small(1.95)}$ $5.52_{\small(1.15)}$
GraphRec $35.68_{\small(2.83)}$ $39.00_{\small(1.68)}$ $26.24_{\small(.01.78)}$ $40.66_{\small(2.71)}$ $40.43_{\small(1.65)}$ $28.85_{\small(2.34)}$ $47.00_{\small(1.81)}$ $42.76_{\small(1.28)}$ $33.26_{\small(1.49)}$
MAMO $60.09_{\small(0.12)}$ $72.78_{\small(0.04)}$ $50.37_{\small(0.11)}$ $67.50_{\small(0.16)}$ $75.31_{\small(0.05)}$ $57.21_{\small(0.16)}$ $\mathbf{79.76_{\small(0.06)}}$ $80.07_{\small(0.03)}$ $\mathbf{71.73_{\small(0.08)}}$
TaNP $60.67_{\small(2.44)}$ $87.34_{\small(0.40)}$ $49.82_{\small(2.17)}$ $\mathbf{69.62_{\small(0.89)}}$ $89.03_{\small(0.30)}$ $58.05_{\small(1.66)}$ $77.55_{\small(1.74)}$ $91.00_{\small(0.39)}$ $66.90_{\small(1.71)}$
MeLU $50.32_{\small(1.73)}$ $67.37_{\small(0.70)}$ $39.34_{\small(1.11)}$ $55.44_{\small(2.73)}$ $68.75_{\small(1.16)}$ $43.55_{\small(2.28)}$ $67.17_{\small(1.25)}$ $72.94_{\small(0.78)}$ $56.26_{\small(1.36)}$
HIRE $\mathbf{62.66_{\small(1.95)}}$ $\mathbf{89.02_{\small(2.14)}}$ $\mathbf{54.16_{\small(5.86)}}$ $69.53_{\small(1.69)}$ $\mathbf{90.17_{\small(1.94)}}$ $\mathbf{61.48_{\small(3.24)}}$ $78.33_{\small(2.51)}$ $\mathbf{92.47_{\small(1.66)}}$ $71.06_{\small(2.00)}$